# Massless Quasiparticles in Bogoliubov-de Gennes Systems

### v1

@@ II.1 Symmetry operations
In the present work, we will not dwell on the structures of all possible fermionic groups $G_{f}$ for a given $G_{b}$  [^1] A.Z. Yang, Z.X. Liu, to appear, we just distinguish two types of them: in the first type, $G_{f}$ is ‘diagonal’ in the particle-hole sector; in the second type (such as projective symmetry groups of $Z_{2}$ QSLs) $G_{f}$ contains symmetry operations that are non-diagonal in the particle-hole sector.
Usually, an eigen space of a Hamiltonian carries an irrep of the symmetry group. However, in BdG systems since the ‘particle-hole symmetry’ $\mathcal{P}$ is anti-commuting with the Hamiltonian, the relationship between the energy eigenstates and the irreps of the group $F_{f}$ is subtle. In the following subsections, we will discuss zero energy modes and nonzero energy modes separately.
II.2 Irreducible Zero modes
Here we systematically study the zero modes in BdG Hamiltonians. We will focus on the spectrum at given momentum $\bm{k}$. We denote the effective particle-hole operation as $\tilde{\mathcal{P}}$ and write the ‘true symmetry group’ as $G_{f}(k)$ which is formed by operations that are commuting with $\mathscr{H}_{k}$. Then the full little co-group $F_{f}(k)$ is formed, with $F_{f}(k)=G_{f}(k)+\tilde{\mathcal{P}}G_{f}(k)$. Suppose that $\mathscr{H}_{k}$ contain zero modes and assume that the zero modes span a Hilbert space $\mathcal{L}_{0}$. We first consider the ‘minimal’ set of zero modes, where $\mathcal{L}_{0}$ carries irreducible Rep $D(F_{f}(k))$ of $F_{f}(k)$.
The restriction of $D(F_{f}(k))$ to the subgroup $G_{f}(k)$ can only have two possibilities (see App.D): it is either an irrep or a direct sum of two irreps of $G_{f}(k)$. If $\mathcal{L}_{0}$ also carries an irrep of $G_{f}(k)$(namely if the restricted rep $D(G_{f}(k))$ is irreducible), then the zero modes are stable and are protected by symmetry. We call this set of zero modes as ‘irreducible’ zero modes. On the other hand, if $\mathcal{L}_{0}$ is a direct sum of two subspaces each carrying an irrep of $G_{f}(k)$, then the zero modes in $\mathcal{L}_{0}$ are accidental and are not symmetry protected. In this case, one can add a small perturbation to lift the zero modes without breaking any symmetry. The obtained eigen modes in $\mathcal{L}_{0}$ have nonzero energies and are called a set of ‘irreducible nonzero modes’ in later discussion.
However, different irreducible zero modes can couple to each other and form nonzero energy modes, which is described by the following theorem.
@@ II.2 Irreducible Zero modes
Every irrep $D(F_{f}(k))$ of a set of zero modes has a partner $D^{\prime}(F_{f}(k))$ with which the zero modes can couple to form nonzero energy modes, if the bases of both $D(F_{f}(k))$ and $D^{\prime}(F_{f}(k))$ are included in the system.
@@
According to the theorem, the ‘protected’ zero modes at the given momentum point $k$ can be lifted if it is allowed to add more orbitals to the system to enlarge the Hilbert space. Consequently, the zero modes are either completely gapped out or moved to a different momentum point.
Suppose that the system contains more than one sets of irreducible zero modes at $k$. If they are not coupleable partners of each other, then these zero modes are symmetry protected and stable. Otherwise, if two sets of irreducible zero modes are coupleable partners of each other, then they can be lifted to finite energy by adiabatically adjusting the Hamiltonian without breaking any symmetry. In the next subsection, we will analyze the precise relation between $D(F_{f}(k))$ and its coupleable partner $D^{\prime}(F_{f}(k))$.
@@ II.3 Minimal nonzero modes: $\{\varepsilon_{k},-\varepsilon_{k}\}$ with $\varepsilon_{k}\neq 0$
Now we discuss the modes having nonzero energy. Since $\tilde{\mathcal{P}}$ maps the $n$-dimensional ($n$-D) eigenspace of $\mathscr{H}_{k}$ with eigenvalue $\varepsilon_{k}$ to the one with eigenvalue $-\varepsilon_{k}$, we need to investigate the $2n$-D subspace ±$\mathcal{L}_{\pm\varepsilon}$ formed by the direct sum of the eigenspaces $\{\varepsilon_{k},-\varepsilon_{k}\}$ for $\varepsilon_{k}\neq 0$. This subspace ±$\mathcal{L}_{\pm\varepsilon}$ is called minimal if it cannot be divided into smaller sets of nonzero modes without breaking any symmetry. There are two possible types of minimal nonzero modes: (i) ±$\mathcal{L}_{\pm\varepsilon}$ carries an irreducible rep of $F_{f}(k)$ and is called a set of ‘irreducible’ nonzero modes (INZM); (ii) ±$\mathcal{L}_{\pm\varepsilon}$ carries a direct sum of two irreducible zero modes of $F_{f}(k)$, and ±$\mathcal{L}_{\pm\varepsilon}$ will be called a set of ‘reducible’ minimal nonzero modes (RMNZM).
@@ II.3.1 Irreducible nonzero modes
By definition, the supporting space ±$\mathcal{L}_{\pm\varepsilon}$ of irreducible nonzero modes forms an irrep $D(F_{f}(k))$. The restricted rep on $G_{f}(k)$, namely $D(G_{f}(k))$, must be reducible, otherwise it would gives rise to a set of irreducible zero modes. We adopt the eigenbases of the Hamiltonian for the irreducible nonzero modes which takes the form $h(k)=\varepsilon_{k}\mu_{z}\otimes I_{n}$, where $\mu_{z}$ is the third Pauli matrix whose eigenvalues ±$\pm 1$ respectively label the eigen spaces with positive and negative energy. In the induced bases, see App.D, the element $g\in G_{f}(k)$ is represented as
$D(g)K_{s(g)}=\begin{pmatrix}M(g)&0\\ 0&M^{\prime}(g)\end{pmatrix}K_{s(g)},$
where $M(g)K_{s(g)}$ and $M^{\prime}(g)K_{s(g)}={\omega_{2}(g,\tilde{\mathcal{P}})\over\omega_{2}(\tilde{\mathcal{P}},\tilde{\mathcal{P}}^{-1}g\tilde{\mathcal{P}})}K_{s(\tilde{\mathcal{P}})}M(\tilde{\mathcal{P}}^{-1}g\tilde{\mathcal{P}})K_{s(g\tilde{\mathcal{P}})}$ are two $n$-D irreps of $G_{f}(k)$, and $\tilde{\mathcal{P}}$ is represented as
$D(\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}=\begin{pmatrix}0&\omega_{2}(\tilde{\mathcal{P}},\tilde{\mathcal{P}})M(\tilde{\mathcal{P}}^{2})\\ I_{n}&0\end{pmatrix}K_{s(\tilde{\mathcal{P})}}.$
@@ II.3.2 Reducible minimal nonzero modes
We now consider the case in which ±$\mathcal{L}_{\pm\varepsilon}$ contains two irreps of $F_{f}(k)$ having the the same factor system, and we note them as $D(F_{f}(k))$ and $D^{\prime}(F_{f}(k))$ respectively. By definition, the ‘minimal nonzero modes’ requires that the restricted Reps $D(G_{f}(k))$ and $D^{\prime}(G_{f}(k))$ are irreducible. From Schur’s lemma, the restricted reps $D(G_{f}(k))$ and $D^{\prime}(G_{f}(k))$ must be equivalent such that they can couple to each other. That is, $D(G_{f}(k))\cong D^{\prime}(G_{f}(k))$. We thus assume that the bases have been adjusted such that
@@
Furthermore, as summarized in table 1, the ability of coupling between the two sets of irreducible zero modes further restricts the relation between $D(\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}$ and $D^{\prime}(\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}$. In later discussion, we define the unitary quantity
$M=D(\tilde{\mathcal{P}})[D^{\prime}(\tilde{\mathcal{P}})]^{-1}$
to denote the difference between $D(\tilde{\mathcal{P}})$ and $D^{\prime}(\tilde{\mathcal{P}})$. As shown in App.E, the $M$ is commuting with the restricted rep $D(G_{f}(k))$.
@@
(a) If $G_{f}(k)$ is unitary, then from Schur’s lemma one has $M=\eta_{\tilde{\mathcal{P}}}I_{n}$ with $\eta_{\tilde{\mathcal{P}}}$ a constant. Since $\tilde{\mathcal{P}}^{2}\in G_{f}(k)$, hence $[D(\tilde{\mathcal{P}})]^{2}=[D^{\prime}(\tilde{\mathcal{P}})]^{2}$, namely $\eta_{\tilde{\mathcal{P}}}^{2}=1$, $\eta_{\tilde{\mathcal{P}}}=\pm 1$, see App.E.1.1. When $\eta_{\tilde{\mathcal{P}}}=-1$, then ±$\mathcal{L}_{\pm\varepsilon}$ is indeed a set of reducible minimal nonzero modes. When $\eta_{\tilde{\mathcal{P}}}=1$, ±$\mathcal{L}_{\pm\varepsilon}$ contains two sets of irreducible zero modes uncoupled to each other.
(b) If $G_{f}(k)$ is anti-unitary, denote the maximum unitary subgroup of $G_{f}(k)$ as $H_{f}(k)$. Then there will be three different situations.
b1), When the restricted rep $D(G_{f}(k))$ is of the real class $\mathbb{R}$, then $M=\eta_{\tilde{\mathcal{P}}}I_{n}$ still holds with $\eta_{\tilde{\mathcal{P}}}=\pm 1$. The case with $\eta_{\tilde{\mathcal{P}}}=-1$ stands for RMNZM.
b2), When the $D(G_{f}(k))$ is of the complex class $\mathbb{C}$, then the restricted rep $D(H_{f}(k))$ is a direct sum of two non-equivalent irreps with dimension ${n\over 2}$. Suppose $\Omega_{z}$ is diagonal in this direct sum space and is represented as $\Omega_{z}=\omega_{z}\otimes I_{n\over 2}$ with $\omega_{z}$ the third pauli matrix. Then the unitary elements in the centralizer of the rep $D(G_{f}(k))$ can be written as $e^{\mathrm{i}\theta\Omega_{z}}=\cos\theta+\mathrm{i}\sin\theta\Omega_{z}$, see App.E.1.2, where $\theta\in[0,2\pi)$. So we have $M=e^{\mathrm{i}\theta_{0}\Omega_{z}}$ for some $\theta_{0}$ with a constraint,
@@
If $\{D(\tilde{\mathcal{P}}),\Omega_{z}\}\neq 0$, then $\sin\theta_{0}=0$, so $\theta_{0}=0,\pi$. The case $\theta_{0}=\pi$ (yielding $M=-I_{n}$) stands for the RMNZM. If $\{D(\tilde{\mathcal{P}}),\Omega_{z}\}=0$, then $\theta_{0}\in[0,2\pi)$ and $D(F_{f}(k))$ and $D^{\prime}(F_{f}(k))$ are equivalent since they are related by a $U(1)$ transformation $\Psi^{\dagger}\to\Psi^{\prime{\dagger}}=\Psi^{\dagger}e^{-\mathrm{i}\Omega_{z}{\theta_{0}\over 2}}$, $D(\tilde{\mathcal{P}})\to D^{\prime}(\tilde{\mathcal{P}})=e^{-\mathrm{i}\Omega_{z}{\theta_{0}\over 2}}D(\tilde{\mathcal{P}})e^{\mathrm{i}\Omega_{z}{\theta_{0}\over 2}}=e^{-\mathrm{i}\Omega_{z}{\theta_{0}}}D(\tilde{\mathcal{P}})$. The two irreps can be coupled to form nonzero modes. For instance, when $\theta_{0}=0$ the Hamiltonian can be chosen as $h_{k}=\varepsilon_{k}\mu_{y}\otimes\Omega_{z}.$
@@
$M=e^{\mathrm{i}\theta_{0}\Omega_{\boldsymbol{n}_{0}}}.$ (11)
@@
$e^{-\mathrm{i}\theta_{0}\Omega_{\boldsymbol{n}_{0}}}D(\tilde{\mathcal{P}})=D(\tilde{\mathcal{P}})e^{\mathrm{i}\theta_{0}\Omega_{\boldsymbol{n}_{0}}}$ (12)
If $\{D(\tilde{\mathcal{P}}),\Omega_{\boldsymbol{n}_{0}}\}\neq 0$ for any ${\boldsymbol{n}_{0}}\in S^{2}$ (for instance, when ${\rm Tr\ D(\tilde{\mathcal{P}})}\neq 0$), then only $\theta_{0}=0,\pi$ satisfies the relation (10). Then (11) reduces to $M=\eta_{\tilde{\mathcal{P}}}I_{n}$ with $\eta_{\tilde{\mathcal{P}}}=\pm 1$ and the case $\eta_{\tilde{\mathcal{P}}}=-1$ stands for RMNZM. If there exist some ${\boldsymbol{n}_{0}}$ such that $\{D(\tilde{\mathcal{P}}),\Omega_{\boldsymbol{n}_{0}}\}=0$, then $D(\tilde{\mathcal{P}})$ and $D^{\prime}(\tilde{\mathcal{P}})$ are related by an SU(2) transformation $\Psi^{\dagger}\to\Psi^{\prime{\dagger}}=\Psi^{\dagger}e^{-\mathrm{i}\Omega_{\boldsymbol{n}_{0}}{\theta_{0}\over 2}}$, $D(\tilde{\mathcal{P}})\to D^{\prime}(\tilde{\mathcal{P}})=e^{-\mathrm{i}\Omega_{\boldsymbol{n}_{0}}{\theta_{0}\over 2}}D(\tilde{\mathcal{P}})e^{\mathrm{i}\Omega_{\boldsymbol{n}_{0}}{\theta_{0}\over 2}}=e^{-\mathrm{i}\Omega_{\boldsymbol{n}_{0}}{\theta_{0}}}D(\tilde{\mathcal{P}})$. Hence, the two irreps $D(F_{f}(k)$ and $D^{\prime}(F_{f}(k))$ are equivalent, they can be coupled to form nonzero modes. For instance, when $\theta_{0}=0$ the Hamiltonian can be chosen as $h_{k}=\varepsilon_{k}\mu_{y}\otimes\Omega_{n}.$
@@
If $G_{f}(k)$ is anti-unitary, we can multiply $\tilde{\mathcal{P}}$ with an anti-unitary element $T_{0}\in G_{f}(k)$ to obtain a chiral-like symmetry operator $\tilde{\mathcal{P}}^{\prime}=\tilde{\mathcal{P}}T_{0}$, which has been discussed in case (I). Therefore, we only need to consider the case in which $G_{f}(k)$ is unitary. Eq. (9) indicates that the two irreps $D(F_{f}(k))$ and $D^{\prime}(F_{f}(k))$ are equivalent, and consequently $D^{\prime}(\tilde{\mathcal{P}})K=e^{-\mathrm{i}\theta}D(\tilde{\mathcal{P}})K$ with $\theta\in[0,2\pi)$, namely $M=e^{\mathrm{i}\theta}I_{n}$, see App.E.2. Now ±$\mathcal{L}_{\pm\varepsilon}$ is a set of RMNZM since there exist a Hamiltonian matrix $h_{k}=\varepsilon_{k}(\sin{\theta\over 2}\mu_{x}+\cos{\theta\over 2}\mu_{y})\otimes I_{n}$ which commutes with $\mathscr{D}(G_{f}(k))=D(G_{f}(k))\oplus D(G_{f}(k))$ and anti-commutes with $\mathscr{D}(\tilde{\mathcal{P}})K=D(\tilde{\mathcal{P}})\oplus e^{-\mathrm{i}\theta}D(\tilde{\mathcal{P}})K$.

### v3

@@ II.1 Symmetry operations
In the present work, we will not dwell on the structures of all possible fermionic groups $G_{f}$ for a given $G_{b}$, and we just distinguish two types of them: in the first type, $G_{f}$ is ‘diagonal’ in the particle-hole sector (such as the double MSGs/SSGs); in the second type, $G_{f}$ contains symmetry operations that are non-diagonal in the particle-hole sector (such as projective symmetry groups of $Z_{2}$ QSLs).
In the following discussion, we will focus on the spectrum at a momentum $\bm{k}$ that is invariant under an effective particle-hole symmetry $\tilde{\mathcal{P}}$, since zero modes at momentum $\bm{k}$ without this symmetry are unstable when varying the chemical potential. Denoting the group of symmetry operations that commute with $\mathscr{H}_{k}$ as $G_{f}(k)$, then the full little co-group $F_{f}(k)$ is given by $F_{f}(k)=G_{f}(k)+\tilde{\mathcal{P}}G_{f}(k)$. Since $[\mathscr{H}_{k},G_{f}(k)]=0$, generally the eigenstates of $\mathscr{H}_{k}$ with eigvalue $\varepsilon_{k}$ carry an irreducible projective representation of $G_{f}(k)$ (the factor systems $\omega_{2}(g_{1},g_{2})$ are given in App.C). Noticing that $\tilde{\mathcal{P}}$ maps $|\varepsilon_{k}\rangle$ to $|-\varepsilon_{k}\rangle$, in order to obtain more information of the energy spectrum especially the existence or nonexistence of zero energy modes, we need to consider the full group $F_{f}(k)$ and its irreps.
$F_{f}(k)$ has two kinds of projective irreps (see App.D):
(1)simple irreps, which are irreducible when restricted to the subgroup $G_{f}(k)$;
(2)composite irreps, the restricted rep of $G_{f}(k)$ for a composite irreps is a direct sum of two irreps of $G_{f}(k)$.
II.2 Robust and accidental Zero modes
Here we systematically study the zero modes in BdG Hamiltonians. Suppose that $\mathscr{H}_{k}$ contain zero modes and assume that the zero modes span a Hilbert space $\mathcal{L}_{0}$.
(i) If $\mathcal{L}_{0}$ carries a simple irrep of $F_{f}(k)$, namely if $\mathcal{L}_{0}$ also carries an irrep of $G_{f}(k)$, then the zero modes are stable in the sense that they cannot be lifted when keeping all of the symmetries. We call this set of zero modes as ‘irreducible zero modes’.
Although a set of irreducible zero modes carry a simple irrep of $F_{f}(k)$, the inverse is not necessarily true. Namely, a simple irrep of $F_{f}(k)$ may not contribute any zero modes to the spectrum because two simple irreps may couple with each other to form nonzero energy modes. So we have the following theorem.
@@ II.2 Robust and accidental Zero modes
Every simple projective irrep $D(F_{f}(k))$ has a partner $D^{\prime}(F_{f}(k))$ (carrying the same factor system of $D(F_{f}(k))$) with whom it can couple to form nonzero energy modes if the bases of $D(F_{f}(k))$ and $D^{\prime}(F_{f}(k))$ are both included in the system.
@@
According to theorem 1, the irrep $D(G_{f}(k))$ gives rise to ‘irreducible zero modes’ only when its partner $D^{\prime}(F_{f}(k))$ is absent in the bases of the system. Later we will show that this is a stringent condition and it happens only when the system contains special symmetries.
(ii) If $\mathcal{L}_{0}$ carries a composite irrep of $F_{f}(k)$ (namely if $\mathcal{L}_{0}$ carries a reducible rep of $G_{f}(k)$), then the zero modes in $\mathcal{L}_{0}$ are accidental and can be lifted by small perturbations without breaking any symmetry.
(iii) If $\mathcal{L}_{0}$ carries a reducible rep of $F_{f}(k)$, then the zero modes contributed by composite irreps or couples of simple irreps (the simple irreps in each couple are partners of each other) are accidental, while the ones contributed by uncoupled simple irreps are robust.
@@ II.3 Minimal nonzero modes: $\{\varepsilon_{k},-\varepsilon_{k}\}$ with $\varepsilon_{k}\neq 0$
Now we discuss the modes having nonzero energy. Since $\tilde{\mathcal{P}}$ maps the $n$-dimensional ($n$-D) eigenspace of $\mathscr{H}_{k}$ with eigenvalue $\varepsilon_{k}$ to the one with eigenvalue $-\varepsilon_{k}$, we need to investigate the $2n$-D subspace ±$\mathcal{L}_{\pm\varepsilon_{k}}$ formed by the direct sum of the eigenspaces $\{\varepsilon_{k},-\varepsilon_{k}\}$ for $\varepsilon_{k}\neq 0$. This subspace ±$\mathcal{L}_{\pm\varepsilon_{k}}$ is called minimal if it cannot be divided into smaller sets of nonzero modes without breaking any symmetry. There are two possible types of minimal nonzero modes: (1) ±$\mathcal{L}_{\pm\varepsilon_{k}}$ carries a composite irreducible rep of $F_{f}(k)$ and is called a set of ‘irreducible nonzero modes’; (2) ±$\mathcal{L}_{\pm\varepsilon_{k}}$ carries a direct sum of two simple irreps of $F_{f}(k)$, and ±$\mathcal{L}_{\pm\varepsilon_{k}}$ will be called a set of ‘reducible minimal nonzero modes’ (RMNZM).
@@ II.3.1 Irreducible nonzero modes
By definition, the supporting space ±$\mathcal{L}_{\pm\varepsilon_{k}}$ of irreducible nonzero modes forms an irrep of $D(F_{f}(k))$. This irrep must be a composite irrep, namely, the restricted rep on $G_{f}(k)$ must be reducible, otherwise it would gives rise to a set of irreducible zero modes. We adopt the eigenbases of the Hamiltonian for the irreducible nonzero modes which takes the form $h(k)=\varepsilon_{k}\mu_{z}\otimes I_{n}$, where $\mu_{z}$ is the third Pauli matrix whose eigenvalues ±$\pm 1$ respectively label the eigen spaces with positive and negative energy. Defining $K_{-1}=K,K_{1}=I$, and $s(g)=-1$ for anti-unitary elements $g\in G_{f}(k)$ and $s(g)=1$ for unitary ones, then according to App.D the element $g\in G_{f}(k)$ is represented as
$D(g)K_{s(g)}=\begin{pmatrix}d(g)&0\\ 0&d^{\prime}(g)\end{pmatrix}K_{s(g)},$
where $d(g)K_{s(g)}$ and $d^{\prime}(g)K_{s(g)}={\omega_{2}(g,\tilde{\mathcal{P}})\over\omega_{2}(\tilde{\mathcal{P}},\tilde{\mathcal{P}}^{-1}g\tilde{\mathcal{P}})}K_{s(\tilde{\mathcal{P}})}d(\tilde{\mathcal{P}}^{-1}g\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}$ are two $n$-D irreps of $G_{f}(k)$, and $\tilde{\mathcal{P}}$ is represented as
$D(\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}=\begin{pmatrix}0&\omega_{2}(\tilde{\mathcal{P}},\tilde{\mathcal{P}})d(\tilde{\mathcal{P}}^{2})\\ I_{n}&0\end{pmatrix}K_{s(\tilde{\mathcal{P})}}$
with $\tilde{\mathcal{P}}^{2}\in G_{f}(k)$.
@@ II.3.2 Reducible minimal nonzero modes
We now consider the case in which ±$\mathcal{L}_{\pm\varepsilon_{k}}$ contains two simple irreps of $F_{f}(k)$ noted as $D(F_{f}(k))$ and $D^{\prime}(F_{f}(k))$ respectively (they are partner of each other). By definition, the ‘minimal nonzero modes’ requires that the restricted Reps $D(G_{f}(k))$ and $D^{\prime}(G_{f}(k))$ are irreducible. From Schur’s lemma, the restricted reps $D(G_{f}(k))$ and $D^{\prime}(G_{f}(k))$ must be equivalent such that they can couple to each other. That is, $D(G_{f}(k))\cong D^{\prime}(G_{f}(k))$. We thus assume that the bases have been adjusted such that
@@
Furthermore, as summarized in table 1, the ability of coupling between the two simple irreps further restricts the relation between $D(\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}$ and $D^{\prime}(\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}$. In later discussion, we define the unitary quantity
$X=D(\tilde{\mathcal{P}})[D^{\prime}(\tilde{\mathcal{P}})]^{-1}$
to denote the difference between $D(\tilde{\mathcal{P}})$ and $D^{\prime}(\tilde{\mathcal{P}})$. As shown in App.E, $X$ is commuting with the restricted rep $D(G_{f}(k))$.
@@
(a) If $G_{f}(k)$ is unitary, then from Schur’s lemma one has $X=\eta_{\tilde{\mathcal{P}}}I_{n}$ with $\eta_{\tilde{\mathcal{P}}}$ a constant. Since $\tilde{\mathcal{P}}^{2}\in G_{f}(k)$, hence $[D(\tilde{\mathcal{P}})]^{2}=[D^{\prime}(\tilde{\mathcal{P}})]^{2}$, namely $\eta_{\tilde{\mathcal{P}}}^{2}=1$, $\eta_{\tilde{\mathcal{P}}}=\pm 1$, see App.E.1.1. When $\eta_{\tilde{\mathcal{P}}}=-1$, then ±$\mathcal{L}_{\pm\varepsilon_{k}}$ is indeed a set of RMNZM. When $\eta_{\tilde{\mathcal{P}}}=1$, ±$\mathcal{L}_{\pm\varepsilon_{k}}$ contains two simple irreps uncoupled to each other, hence ±$\mathcal{L}_{\pm\varepsilon_{k}}$ corresponds to zero modes.
(b) If $G_{f}(k)$ is anti-unitary, denote the maximum unitary subgroup of $G_{f}(k)$ as $H_{f}(k)$ with $G_{f}(k)=H_{f}(k)+T_{0}H_{f}(k),\ T_{0}^{2}\in H_{f}(k)$. The projective irreps of $G_{f}(k)$ are classified into $\mathbb{R}$, $\mathbb{C}$, and $\mathbb{H}$ classes 111 The class is determined by the Wigner indicator of an irrep $D(G_{f}(k))$ of $G_{f}(k)=H_{f}(k)+T_{0}H_{f}(k)$, namely $W_{D}={1\over|H_{f}(k)|}\sum_{h\in H_{f}(k)}\omega_{2}(T_{0}h,T_{0}h){\rm Tr}\left[D(T_{0}h)^{2}\right]$. The values of the Wigner indicator in the $\mathbb{R}$, $\mathbb{C}$ and $\mathbb{H}$ classes are $W_{D}^{\mathbb{R}}=1,W_{D}^{\mathbb{C}}=0,W_{D}^{\mathbb{H}}=-2$ respectively. according to their centralizersShaw and Lever [1974].
b1), When the restricted rep $D(G_{f}(k))$ is of the real class $\mathbb{R}$, then $X=\eta_{\tilde{\mathcal{P}}}I_{n}$ still holds with $\eta_{\tilde{\mathcal{P}}}=\pm 1$. The case with $\eta_{\tilde{\mathcal{P}}}=-1$ stands for RMNZM.
b2), When the $D(G_{f}(k))$ is of the complex class $\mathbb{C}$, then the restricted rep $D(H_{f}(k))$ is a direct sum of two non-equivalent irreps with dimension ${n\over 2}$. Suppose $\Omega_{z}$ is diagonal in this direct sum space and is represented as $\Omega_{z}=\omega_{z}\otimes I_{n\over 2}$ with $\omega_{z}$ the third pauli matrix. Then the unitary elements in the centralizer of the rep $D(G_{f}(k))$ can be written as $e^{\mathrm{i}\theta\Omega_{z}}=\cos\theta+\mathrm{i}\sin\theta\Omega_{z}$, see App.E.1.2, where $\theta\in[0,2\pi)$. So we have $X=e^{\mathrm{i}\theta_{0}\Omega_{z}}$ for some $\theta_{0}$ with a constraint,
@@
If $\{D(\tilde{\mathcal{P}}),\Omega_{z}\}\neq 0$, then $\sin\theta_{0}=0$, so $\theta_{0}=0,\pi$. The case $\theta_{0}=\pi$ (yielding $X=-I_{n}$) stands for the RMNZM. If $\{D(\tilde{\mathcal{P}}),\Omega_{z}\}=0$, then $\theta_{0}\in[0,2\pi)$ and $D(F_{f}(k))$ and $D^{\prime}(F_{f}(k))$ are equivalent since they are related by a $U(1)$ transformation $\Psi^{\dagger}\to\Psi^{\prime{\dagger}}=\Psi^{\dagger}e^{-\mathrm{i}\Omega_{z}{\theta_{0}\over 2}}$, $D(\tilde{\mathcal{P}})\to D^{\prime}(\tilde{\mathcal{P}})=e^{-\mathrm{i}\Omega_{z}{\theta_{0}\over 2}}D(\tilde{\mathcal{P}})e^{\mathrm{i}\Omega_{z}{\theta_{0}\over 2}}=e^{-\mathrm{i}\Omega_{z}{\theta_{0}}}D(\tilde{\mathcal{P}})$. The two irreps can be coupled to form nonzero modes. For instance, when $\theta_{0}=0$ the Hamiltonian can be chosen as $h_{k}=\varepsilon_{k}\mu_{y}\otimes\Omega_{z}.$
@@
$X=e^{\mathrm{i}\theta_{0}\Omega_{\boldsymbol{n}_{0}}}.$ (11)
@@
$e^{-\mathrm{i}\theta_{0}\Omega_{\boldsymbol{n}_{0}}}D(\tilde{\mathcal{P}})=D(\tilde{\mathcal{P}})e^{\mathrm{i}\theta_{0}\Omega_{\boldsymbol{n}_{0}}}.$ (12)
If $\{D(\tilde{\mathcal{P}}),\Omega_{\boldsymbol{n}_{0}}\}\neq 0$ for any ${\boldsymbol{n}_{0}}\in S^{2}$ (for instance, when ${\rm Tr\ D(\tilde{\mathcal{P}})}\neq 0$), then only $\theta_{0}=0,\pi$ satisfies the relation (10). Then (11) reduces to $X=\eta_{\tilde{\mathcal{P}}}I_{n}$ with $\eta_{\tilde{\mathcal{P}}}=\pm 1$ and the case $\eta_{\tilde{\mathcal{P}}}=-1$ stands for RMNZM. If there exist some ${\boldsymbol{n}_{0}}$ such that $\{D(\tilde{\mathcal{P}}),\Omega_{\boldsymbol{n}_{0}}\}=0$, then $D(\tilde{\mathcal{P}})$ and $D^{\prime}(\tilde{\mathcal{P}})$ are related by an SU(2) transformation $\Psi^{\dagger}\to\Psi^{\prime{\dagger}}=\Psi^{\dagger}e^{-\mathrm{i}\Omega_{\boldsymbol{n}_{0}}{\theta_{0}\over 2}}$, $D(\tilde{\mathcal{P}})\to D^{\prime}(\tilde{\mathcal{P}})=e^{-\mathrm{i}\Omega_{\boldsymbol{n}_{0}}{\theta_{0}\over 2}}D(\tilde{\mathcal{P}})e^{\mathrm{i}\Omega_{\boldsymbol{n}_{0}}{\theta_{0}\over 2}}=e^{-\mathrm{i}\Omega_{\boldsymbol{n}_{0}}{\theta_{0}}}D(\tilde{\mathcal{P}})$. Hence, the two irreps $D(F_{f}(k)$ and $D^{\prime}(F_{f}(k))$ are equivalent, they can be coupled to form nonzero modes. For instance, when $\theta_{0}=0$ the Hamiltonian can be chosen as $h_{k}=\varepsilon_{k}\mu_{y}\otimes\Omega_{n}.$
@@
If $G_{f}(k)$ is anti-unitary, we can multiply $\tilde{\mathcal{P}}$ with an anti-unitary element $T_{0}\in G_{f}(k)$ to obtain a chiral-like symmetry operator $\tilde{\mathcal{P}}^{\prime}=\tilde{\mathcal{P}}T_{0}$, which has been discussed in case (I). Therefore, we only need to consider the case in which $G_{f}(k)$ is unitary. Eq. (9) indicates that the two irreps $D(F_{f}(k))$ and $D^{\prime}(F_{f}(k))$ are equivalent, and consequently $D^{\prime}(\tilde{\mathcal{P}})K=e^{-\mathrm{i}\theta}D(\tilde{\mathcal{P}})K$ with $\theta\in[0,2\pi)$, namely $X=e^{\mathrm{i}\theta}I_{n}$, see App.E.2. Now ±$\mathcal{L}_{\pm\varepsilon}$ is a set of RMNZM since there exist a Hamiltonian matrix $h_{k}=\varepsilon_{k}(\sin{\theta\over 2}\mu_{x}+\cos{\theta\over 2}\mu_{y})\otimes I_{n}$ which commutes with $\mathscr{D}(G_{f}(k))=D(G_{f}(k))\oplus D(G_{f}(k))$ and anti-commutes with $\mathscr{D}(\tilde{\mathcal{P}})K=D(\tilde{\mathcal{P}})\oplus e^{-\mathrm{i}\theta}D(\tilde{\mathcal{P}})K$.